OpenStudy (anonymous):

A 1975 kg car is parked at the top of a steep 42 m long hill inclined at an angle of 15 degrees. If the car starts rolling down the hill, how fast will it be going when it reaches the bottom of the hill? (Neglect friction)

OpenStudy (anonymous):

Assuming no friction, massless tire, ooor let's assume it's just a block on frictionless surface :)  \[a_x=\frac {F_x}{m}\\a_x=\frac{mg\sin15}{m}\\a_x=g \sin15\\v^2=v_0^2+2ax\\v^2=0+2g\sin15*42 \]